Grade 12th passPhysical Chemistry

What is the normality of 5 volumes of H2O2. please give me the answer as soon as possible

Jyotiprakash
9 Years agoGrade 12th pass

3 Answers

Subrahmanyam Kancharla
9 Years ago
Volume strength =5Normality=?Volume strength =5.6*NormalityNormality=Volume strength/Normality =5/5.6 =0.892N
Roma
9 Years ago
1 Normal of H2O2 give 5.6 L oF O2 At STP , so it's volume strength is 5.6
So5.6 V has 1 N
Then 5V will be = 1/5.6 ×5 =5/5.6 = 0.892 N
Uditman
8 Years ago
5V H2O2, it means 1L of H2O2 solution will produce 5L oxygen.22.4L of O2 will produce 68gm H2O2. 5L of O2 will produce by 68*5/22.4 =15.18gm H2O2 Equivalent mass of H2O2=molar mass/2 =32/2=17gmNormality =mass/ (equivalent mass* volume) =15.18/17*1=0.892N